# Lagrangian Motion Fields for Long-term Motion Generation

### v1

@@ 3.1 Lagrangian Motion Fields
Human motion is inherently dynamic with movements unfolding over time in a coordinated manner, featuring natural flow and coherence. Previous work either treats motion as isolated spatial positions, neglecting temporal domain information explicitly [9, 5, 13], or relies on latent-based representations [4, 24, 25], which are difficult to interpret and generalize. To better capture the continuous nature of human motion over time, we propose to apply Lagrangian motion fields for 3D human motion generation.
The Lagrangian perspective, traditionally used in fluid mechanics, tracks individual particles as they move through space and time, rather than using static spatial reference points. Given the initial position $x_{0}$ of a particle, its position at any later time $t$ is determined by the Lagrangian mapping:
@@
Building on this concept, we analogously treat each joint as a particle, whose trajectory is decided by time and the control signal from an initial position. To borrow this concept into the context of general 3D human motion, we further postulate the following assumptions:
  • Continuity of Motion: It is assumed that the positions and velocities of the joints are continuous without any abrupt changes or discontinuities.
  • Static Environment: The environment is assumed to be static, and the motion of the human body is the only dynamic element.
Under these assumptions, the kinematical behavior of each joint over time can be similarly quantified and predicted using Lagrangian motion fields as follows:
@@
where $x_{i}^{j}$ is the representation of the joint $j$ at frame $i$, $\mathbf{C}$ is the control signal and $\mathbf{L}$ is the Lagrangian motion fields function. Notably, the proposed pipeline is not constrained to a specific coordinate system and can be applied universally. Through this formulation, our key insight is to shift from predicting discrete static spatial information snapshots at each frame (i.e., the framewise representation) to modeling the continuous flow of motion.
@@ 3.2 Supermotion
The primary challenge in long-term motion generation is representing lengthy motion sequences compactly while retaining essential information. Drawing inspiration from superpixel techniques in 2D image processing [11, 12, 28], we propose Lagrangian-based supermotion as a condensed representation of prolonged motion sequences, capturing fundamental motion attributes. Each supermotion encapsulates a collection of motion motifs with similar characteristics, thereby reducing the temporal resolution of the sequence and simplifying the generation process.
As illustrated in Fig. 3, for a lengthy motion sequence $\mathbf{X}$, we initially compute the Lagrangian motion flow across the entire sequence. Subsequently, we apply a K-means clustering $\phi(\cdot)$ on the motion fields to classify each part of the motion sequence. After obtaining the motion labels, we smooth the labels and group the adjacent identical labels to form coherent segments $\{\mathbf{seg}_{s}\}_{s=0}^{M-1}$.
@@ 3.4 Refinement Module
where $\beta_{\tau}$ denotes the pre-defined noise variance schedule, []delimited-[][$\cdot$][ ⋅ ] denotes concatenation. To avoid any confusion, we use $\tau$ to denotes the diffusion denoising timestep, distinguishing it from $t$, which represents the frame index.
@@ 4 Experiment
4.1 Long-term Music-to-Dance
Datasets. We conduct the long-term music-to-dance evaluation using the FineDance dataset [43], which comprises 7.7 hours of 30 FPS music and dance motion pairs, covering 16 dance genres, totaling 831,600 frames.
The average duration of each dance segment is 152.3 seconds. In contrast to the commonly used AIST++ dataset [1], which has an average duration of only 13.3 seconds, FineDance better aligns with our long-term setting and poses a greater challenge. During preprocessing, we employ mini-batch K-Means with 1,000 clusters.
4.1.1 Implementation details
4.1.2 Quantitative Evaluation
@@
Following previous work, we assess the motion quality from four different perspectives: Beat Alignment Score (BAS) [1, 42, 4, 9, 5], Foot Skating Ratio (FSR) [45], and Fréchet Inception Distance (FID) of kinetic ($\mathrm{FID}_{k}$) and geometric ($\mathrm{FID}_{g}$) following [5]. BAS measures the synchronization between the generated dance and the given music. FSR evaluates the proportion of foot sliding in the generated motion, reflecting physical realism. $\mathrm{FID}_{k}$ assesses the physical realism of the motion through speed and acceleration features. $\mathrm{FID}_{g}$ measures the overall choreography quality based on predefined geometric motion templates.
For assessing the diversity, we measure the average feature distance of kinetic ($\mathrm{Div}_{k}$) and geometric ($\mathrm{Div}_{g}$) features extracted by fairmotion [46] following the previous works [4].
Efficiency is evaluated by measuring the average runtime required to generate a dance sequence of 1024 frames.
To further demonstrate the robustness of the proposed method, we provide quantitative results for two most commonly used coordinate systems: Cartesian coordinates (denoted as “3D”) and 6-DOF rotation coordinates (denoted as “6D”). These results offer a more thorough comparison, which we will discuss more in Sec. 4.4.
The quantitative results, detailed in Table I, demonstrate significant improvements achieved by our method.
@@
Notably, our approach shows substantial advancements in $\mathrm{FID}_{g}$ scores, with “Lodge (DDIM) + Ours (6D)” achieving the highest score of 23.39, reflecting superior motion quality. Additionally, “Lodge (DDIM) + Ours (6D)” records the lowest Foot Skating Ratio of 1.45, underscoring the model’s capability to produce natural and physically plausible motions. Furthermore, “EDGE + Ours (3D)” achieves the highest Beat Alignment Score (BAS) of 0.235, demonstrating superior synchronization with music.
@@
Our method demonstrates improvements in motion diversity, particularly in the geometric feature space, as we got higher $\mathrm{Div}_{g}$ scores across multiple configurations compared to baseline methods.
Most importantly, our method drastically reduces the runtime due to the superior performance of our proposed supermotion representation, achieving a maximum acceleration of 71.13% based on EDGE and 44.64% based on Lodge.
TABLE II: Quantitative results on the HumanML3D test set. All methods use the real motion length from the ground truth. $\uparrow$’ indicates the higher the better. $\downarrow$’ indicates the lower the better.$\rightarrow$’ indicates that results are better when the metric is closer to the real distribution. The evaluations were performed 20 times, with $\pm$ indicating the 95% confidence interval. Bold indicates the best result, and underline indicates it surpasses its baseline.
[…]
Ground Truth $0.797^{\pm.002}$ $0.002^{\pm.000}$ $2.974^{\pm.008}$ $9.503^{\pm.065}$
JL2P $0.486^{\pm.002}$ $11.02^{\pm.046}$ $5.296^{\pm.008}$ $7.676^{\pm.058}$
Text2Gesture $0.345^{\pm.002}$ $7.664^{\pm.030}$ $6.030^{\pm.008}$ $6.409^{\pm.071}$
T2M ${0.740^{\pm.003}}$ $1.067^{\pm.002}$ $\textbf{3.340}^{\pm.008}$ $9.188^{\pm.002}$
[…]
Top-1 Top-2 Top-3
Ground Truth $0.321^{\pm 0.005}$ $0.468^{\pm 0.005}$ $0.565^{\pm 0.006}$ $0.002^{\pm 0.001}$ $4.204^{\pm 0.024}$ $6.618^{\pm 0.011}$
MDM $0.155^{\pm 0.002}$ $0.260^{\pm 0.003}$ $0.340^{\pm 0.002}$ $2.315^{\pm 0.110}$ $5.033^{\pm 0.018}$ $4.776^{\pm 0.047}$
MDM + Ours (3D) $0.201^{\pm 0.003}$ $0.338^{\pm 0.004}$ $0.442^{\pm 0.005}$ $1.668^{\pm 0.027}$ $\textbf{4.610}^{\pm 0.019}$ $\textbf{5.842}^{\pm 0.039}$
MDM + Ours (6D) $\textbf{0.213}^{\pm 0.005}$ $\textbf{0.351}^{\pm 0.006}$ $\textbf{0.455}^{\pm 0.007}$ $\textbf{1.626}^{\pm 0.031}$ $4.720^{\pm 0.023}$ $5.537^{\pm 0.048}$
T2M-GPT $0.170^{\pm 0.002}$ $0.286^{\pm 0.003}$ $0.374^{\pm 0.002}$ $2.083^{\pm 0.030}$ $4.800^{\pm 0.027}$ $5.415^{\pm 0.056}$
T2M-GPT + Ours (3D) $0.204^{\pm 0.004}$ $0.343^{\pm 0.004}$ $0.453^{\pm 0.005}$ $1.618^{\pm 0.027}$ $\textbf{4.650}^{\pm 0.024}$ $\textbf{5.774}^{\pm 0.061}$
T2M-GPT + Ours (6D) $\textbf{0.217}^{\pm 0.004}$ $\textbf{0.359}^{\pm 0.006}$ $\textbf{0.470}^{\pm 0.005}$ $\textbf{1.586}^{\pm 0.021}$ $4.550^{\pm 0.020}$ $5.632^{\pm 0.045}$
@@ 4.2.3 Quantitative Evaluation
We evaluate the generated motions based on quality, accuracy, diversity, and efficiency, adhering to the evaluation protocol outlined in MDM [13].
For measuring motion quality, we employ FID [51], R-Precision, and Multimodal Distance [49]. R-Precision measures the accuracy of motion-to-text retrieval by ranking the Euclidean distances between motion and text embeddings, with results reported for Top-1, Top-2, and Top-3 accuracy. FID evaluates the distribution distance between generated and real motions by comparing the extracted motion features. Multimodal Distance calculates the average Euclidean distance between text features and their corresponding generated motion features, assessing the alignment between text and motion.
To assess motion diversity, we assess the variation within a set of generated motions by computing the average Euclidean distance between randomly sampled pairs of motion features, following  [51].
The runtime metric measures the time required to sample one motion instance. For a more comprehensive analysis, we conduct comparisons on both short-term and long-term text-to-motion benchmarks, as presented in Table II and Table III, respectively.
@@
Notably, our method demonstrates a more pronounced advantage in long-term text-to-motion tasks. As detailed in  Table III, our methods outperform its baseline models in both R-Precision and Multimodal Distance on the HumanML3D-MP dataset. Specifically, FID decreases by 28% compared to MDM and by 24% compared to T2M-GPT, indicating a significant improvement in motion quality.
In addition, our method shows improvements in motion diversity, increasing by 22% compared to MDM and by 7% compared to T2M-GPT.
The compact nature of our supermotion representation also contributes to a 75% reduction in runtime compared to the original MDM on HumanML3D. Moreover, while T2M-GPT incurs additional training costs for VQVAE, our supermotion compression model imposes negligible costs, further highlighting the efficiency of our approach.
These comprehensive improvements across quality, accuracy, and diversity metrics confirm the effectiveness of our approach, further establishing its superiority over the baseline methods in generating realistic and diverse long-term motions.
@@
LodgeBailandoEDGEFACTT2M-GPTMDM00%$10$%$20$%$30$%$40$%$50$%$60$%$70$%$80$%$90$%$100$%0.730.730.730.730.860.860.860.860.790.790.790.790.670.670.670.670.80.80.80.80.820.820.820.82Win RateUser Study
@@
4.3 Qualitative Evaluation and User Study
Qualitative comparisons for the text-to-motion and music-to-dance tasks are shown in  Fig. 5 and  Fig. 6, respectively. Additional results can be found in the video demonstrations included in the supplemental materials. Our method is capable of generating more diverse motions over long durations.

### v2

@@ 3.1 Lagrangian Motion Fields
Previous work either treats motion as isolated spatial positions, neglecting explicit temporal domain information [9, 5, 13], or relies on latent-based representations [4, 24, 25], which are difficult to interpret and generalize. On the other hand, human motion, much like fluid flow, follows continuous trajectories and exhibits coordinated dynamics that evolve smoothly over time. This fundamental similarity to fluid dynamics provides a theoretical basis for employing field-based modeling in human movement analysis. Since traditional Lagrangian fields in fluid mechanics are continuous, our approach adapts this concept by employing a discretized approximation for computational efficiency, analogous to finite element methods in computational fluid dynamics. Specifically, human motion maintains consistent semantic meaning over short time intervals, a property rooted in biomechanics. From a biomechanical perspective, this consistency arises from the smooth velocity transitions and continuous trajectories of individual joints. Consequently, modeling human motion as a continuous velocity field naturally captures both its spatial structure and temporal evolution. To effectively model the temporal continuity of human motion, we propose Lagrangian motion fields for 3D human motion generation.
The Lagrangian perspective in fluid dynamics follows individual particles as they move through space and time, rather than relying on fixed spatial reference points, as in the Eulerian framework. Given the initial position $x_{0}$ of a particle, its position at any later time $t$ is determined by the Lagrangian mapping:
@@
Building on this concept, we analogously model each joint as a particle whose trajectory is determined by time and control signals from an initial position, under the assumption that joint positions and velocities remain continuous, without abrupt changes or discontinuities. Given each joint $j$ as a Lagrangian particle, the velocity field is defined as:
${v}(\mathbf{X}_{j},t)=\frac{\mathbf{x}_{j}(t+\Delta t)-\mathbf{x}_{j}(t)}{\Delta t},\quad t\in[t_{s},t_{s}+d_{s}],$ (2)
where $t_{s}$ denotes the start time of a motion segment $s$, and $d_{s}$ is its duration. This constitutes a discrete velocity field defined over joints and time segments. Within each motion segment $s$, the velocity $v(\cdot)$ serves as a local field, governing the motion of all joints during $t\in[t_{s},t_{s}+d_{s}]$. The segment duration $d_{s}$ functions as a Kolmogorov scale [pope2001turbulent], smoothing fine-scale variations for computational tractability while preserving essential motion characteristics. Using this discrete velocity field representation, the kinematical behavior of each joint over time can be quantified and predicted as follows:
@@
where ${x}_{i}^{j}$ is the representation of the joint $j$ at frame $i$, $\mathbf{C}$ is the control signal, and $\mathbf{L}$ is the Lagrangian motion fields function that integrates the velocity field to obtain joint positions.
Notably, the proposed pipeline is not constrained to a specific coordinate system and can be applied universally. Through this formulation, our key insight is to shift from predicting discrete static spatial information snapshots at each frame (i.e., the framewise representation) to modeling the continuous flow of motion.
@@ 3.2 Supermotion
Algorithm 1 Motion to Supermotion Conversion
1:Input: Motion sequence set $\mathcal{M}=\{M_{1},M_{2},...,M_{N}\}$, window size $w$, number of clusters $K$
2:Output: Supermotion sequences $\mathcal{S}$
3: // Phase 1: Compute velocities and assign labels
4:for each motion $M_{i}$ in $\mathcal{M}$ do
5:  $V_{i}[t]\leftarrow M_{i}[t+1]-M_{i}[t],\forall t\in[0,L_{i}-1)$
6:end for
[…]
9: // Phase 2: Smooth labels and find segments
10:for each sequence of labels $l_{i}$ do
11:  $\hat{l}_{i}[t]\leftarrow\mathbf{argmax_{k}}|\{l_{i}[j]=k:j\in[t-w/2,t+w/2]\}|$
12:  $P_{i}\leftarrow\{t:\hat{l}_{i}[t]\neq\hat{l}_{i}[t-1]\}\cup\{0,L_{i}\}$
13:end for
14: // Phase 3: Generate supermotions
15:for each motion $M_{i}$ and points $P_{i}$ do
16:  $S_{i}\leftarrow\emptyset$
17:  for each segment $p_{j},p_{j+1}]$ in $P_{i}$ do
18:   $x_{s}\leftarrow M_{i}[p_{j}]$
19:   $v_{s}\leftarrow$ Mean($V_{i}[p_{j}:p_{j+1}]$)
20:   $d_{s}\leftarrow p_{j+1}-p_{j}$
21:   $S_{i}\leftarrow S_{i}\cup\{(x_{s},v_{s},d_{s})\}$
22:  end for
23:  $\mathcal{S}\leftarrow\mathcal{S}\cup\{S_{i}\}$
24:end for
25:return $\mathcal{S}$
The primary challenge in long-term motion generation is efficiently representing extended motion sequences while preserving essential information. While Lagrangian motion fields provide a structured theoretical framework for capturing motion dynamics in a discrete form, a practical mechanism is needed to obtain meaningful discretized motion segments. Drawing inspiration from superpixel techniques in 2D image processing [11, 12, 28], we introduce supermotion as a fundamental abstract motion unit within our Lagrangian framework. Supermotion segments motion into compact units with similar characteristics, enabling efficient compression while maintaining structural coherence. By encapsulating recurrent motion motifs, supermotion reduces the sequence’s temporal resolution, simplifying the generation process while preserving essential dynamics.
As illustrated in Fig. 3 and detailed in Algorithm 1, for a lengthy motion sequence $\mathbf{X}$, we initially compute the Lagrangian motion flow across the entire sequence. Subsequently, we apply a K-means clustering $\phi(\cdot)$ on the motion fields to classify each part of the motion sequence. After obtaining the motion labels, we smooth the labels and group the adjacent identical labels to form coherent segments $\{\mathbf{seg}_{s}\}_{s=0}^{M-1}$.
@@ 3.4 Refinement Module
where $\beta_{\tau}$ denotes the pre-defined noise variance schedule, $\cdot$] denotes concatenation. To avoid any confusion, we use $\tau$ to denotes the diffusion denoising timestep, distinguishing it from $t$, which represents the frame index.
@@ 4 Experiment
4.1 Long-term Music-to-Dance
Datasets. We conduct the long-term music-to-dance evaluation using the FineDance dataset [43], which comprises 7.7 hours of 30 FPS music and dance motion pairs, covering 16 dance genres, totaling 831,600 frames. The average duration of each dance segment is 152.3 seconds. In contrast to the commonly used AIST++ dataset [1], which has an average duration of only 13.3 seconds, FineDance better aligns with our long-term setting and poses a greater challenge. During preprocessing, we employ mini-batch K-Means with 1,000 clusters.
4.1.1 Implementation details
4.1.2 Quantitative Evaluation
@@
Following previous work, we assess the motion quality from four different perspectives: Beat Alignment Score (BAS) [1, 42, 4, 9, 5], Foot Skating Ratio (FSR) [45], and Fréchet Inception Distance (FID) of kinetic ($\mathrm{FID}_{k}$) and geometric ($\mathrm{FID}_{g}$) following [5]. BAS measures the synchronization between the generated dance and the given music. FSR evaluates the proportion of foot sliding in the generated motion, reflecting physical realism. $\mathrm{FID}_{k}$ assesses the physical realism of the motion through speed and acceleration features. $\mathrm{FID}_{g}$ measures the overall choreography quality based on predefined geometric motion templates. For assessing the diversity, we measure the average feature distance of kinetic ($\mathrm{Div}_{k}$) and geometric ($\mathrm{Div}_{g}$) features extracted by fairmotion [46] following the previous works [4]. Efficiency is evaluated by measuring the average runtime required to generate a dance sequence of 1024 frames.
@@
To further demonstrate the robustness of the proposed method, we provide quantitative results for two most commonly used coordinate systems: Cartesian coordinates (denoted as “3D”) and 6-DOF rotation coordinates (denoted as “6D”). These results offer a more thorough comparison, which we will discuss more in Sec. 4.5. The quantitative results, detailed in Table I, demonstrate significant improvements achieved by our method. Notably, our approach shows substantial advancements in $\mathrm{FID}_{g}$ scores, with “Lodge (DDIM) + Ours (6D)” achieving the highest score of 23.39, reflecting superior motion quality. Additionally, “Lodge (DDIM) + Ours (6D)” records the lowest Foot Skating Ratio of 1.45, underscoring the model’s capability to produce natural and physically plausible motions. Furthermore, “EDGE + Ours (3D)” achieves the highest Beat Alignment Score (BAS) of 0.235, demonstrating superior synchronization with music.
@@
Our method demonstrates improvements in motion diversity, particularly in the geometric feature space, as we got higher $\mathrm{Div}_{g}$ scores across multiple configurations compared to baseline methods. Most importantly, our method drastically reduces the runtime due to the superior performance of our proposed supermotion representation, achieving a maximum acceleration of 71.13% based on EDGE and 44.64% based on Lodge.
4.1.3 Comparison with Learned Compact Representations
Bailando [4] is a VQ-VAE-based motion generation method that encodes motion into discrete latent codes for reconstruction. This approach can be extended for long-term representation learning by applying K-Means clustering to latent codes. To evaluate this, we implement a variant, “Long-term Bailando+K-Means”, as shown in Table I, where we cluster the re-trained encoder outputs of Bailando for codebook initialization. This follows the VQ-VAE training practices in [lancucki2020robust], incorporating proper codebook initialization and training stability optimizations.
However, directly clustering these latent codes fails to account for motion dynamics, velocity continuity, and biomechanical constraints. As shown in Table I, our method outperforms “Long-term Bailando+K-Means” across key metrics, demonstrating superior long-term motion generation. While VQ-VAE effectively reconstructs training samples, our supermotion representation better preserves motion coherence by structuring motion based on velocity patterns rather than discrete latent codes.
@@ 4.2.2 Implementation Details
[…]
Ground Truth $797^{\pm.002} 0.0$$002^{\pm.000}$ $2.974^{\pm.008}$ $9.503^{\pm.065}$ -
JL2P $0.486^{\pm.002}$ $11.02^{\pm.046}$ $5.296^{\pm.008}$ $7.676^{\pm.058}$ -
Text2Gesture $0.345^{\pm.002}$ $7.664^{\pm.030}$ $6.030^{\pm.008}$ $6.409^{\pm.071}$ -
T2M $0.740^{\pm.003}$} $1.067^{\pm.002}$ $\textbf{3.340}^{\pm.008}$ $9.188^{\pm.002}$ -
[…]
Top-1 Top-2 Top-3
Ground Truth $0.321^{\pm 0.005}$ $0.468^{\pm 0.005}$ $565^{\pm 0.006} 0.0$$002^{\pm 0.001}$ $4.204^{\pm 0.024}$ $6.618^{\pm 0.011}$ -
MDM $0.155^{\pm 0.002}$ $0.260^{\pm 0.003}$ $0.340^{\pm 0.002}$ $2.315^{\pm 0.110}$ $5.033^{\pm 0.018}$ $4.776^{\pm 0.047}$ 218.34s
MDM + Ours (3D) $0.201^{\pm 0.003}$ $0.338^{\pm 0.004}$ $0.442^{\pm 0.005}$ $1.668^{\pm 0.027}$ $\textbf{4.610}^{\pm 0.019}$ $\textbf{5.842}^{\pm 0.039}$ 11.83s
MDM + Ours (6D) $\textbf{0.213}^{\pm 0.005}$ $\textbf{0.351}^{\pm 0.006}$ $\textbf{0.455}^{\pm 0.007}$ $\textbf{1.626}^{\pm 0.031}$ $4.720^{\pm 0.023}$ $5.537^{\pm 0.048}$ 11.66s
T2M-GPT $0.170^{\pm 0.002}$ $0.286^{\pm 0.003}$ $0.374^{\pm 0.002}$ $2.083^{\pm 0.030}$ $4.800^{\pm 0.027}$ $5.415^{\pm 0.056}$ 2.30s
T2M-GPT + Ours (3D) $0.204^{\pm 0.004}$ $0.343^{\pm 0.004}$ $0.453^{\pm 0.005}$ $1.618^{\pm 0.027}$ $\textbf{4.650}^{\pm 0.024}$ $\textbf{5.774}^{\pm 0.061}$ 0.54s
T2M-GPT + Ours (6D) $\textbf{0.217}^{\pm 0.004}$ $\textbf{0.359}^{\pm 0.006}$ $\textbf{0.470}^{\pm 0.005}$ $\textbf{1.586}^{\pm 0.021}$ $4.550^{\pm 0.020}$ $5.632^{\pm 0.045}$ 0.53s
@@ 4.2.3 Quantitative Evaluation
We evaluate the generated motions based on quality, accuracy, diversity, and efficiency, adhering to the evaluation protocol outlined in MDM [13]. For measuring motion quality, we employ FID [51], R-Precision, and Multimodal Distance [49]. R-Precision measures the accuracy of motion-to-text retrieval by ranking the Euclidean distances between motion and text embeddings, with results reported for Top-1, Top-2, and Top-3 accuracy. FID evaluates the distribution distance between generated and real motions by comparing the extracted motion features. Multimodal Distance calculates the average Euclidean distance between text features and their corresponding generated motion features, assessing the alignment between text and motion. To assess motion diversity, we assess the variation within a set of generated motions by computing the average Euclidean distance between randomly sampled pairs of motion features, following  [51]. The runtime metric measures the time required to sample one motion instance. For a more comprehensive analysis, we conduct comparisons on both short-term and long-term text-to-motion benchmarks, as presented in Table II and Table III, respectively.
@@
Notably, our method demonstrates a more pronounced advantage in long-term text-to-motion tasks. As detailed in  Table III, our methods outperform its baseline models in both R-Precision and Multimodal Distance on the HumanML3D-MP dataset. Specifically, FID decreases by 28% compared to MDM and by 24% compared to T2M-GPT, indicating a significant improvement in motion quality. In addition, our method shows improvements in motion diversity, increasing by 22% compared to MDM and by 7% compared to T2M-GPT.
The compact nature of our supermotion representation also contributes to a 75% reduction in runtime compared to the original MDM on HumanML3D. Moreover, while T2M-GPT incurs additional training costs for VQVAE, our supermotion compression model imposes negligible costs, further highlighting the efficiency of our approach. These comprehensive improvements across quality, accuracy, and diversity metrics confirm the effectiveness of our approach, further establishing its superiority over the baseline methods in generating realistic and diverse long-term motions.
Figure 5: Qualitative visualization of our method across different motion types: (a) Dance movements; (b) Multi-phase sequential actions; (c) Cyclic locomotion; (d) Large-amplitude steady motions.
Figure 6: Text-supermotion alignment visualization. Each example presents the motion sequence (top) and the corresponding attention heatmap (bottom), illustrating the attention weights between text tokens and temporal segments. Darker colors indicate stronger attention weights.
@@
LodgeBailandoEDGEFACTT2M-GPTMDM0%1010%2020%3030%4040%5050%6060%7070%8080%9090%100100%0.730.730.860.860.790.790.670.670.80.80.820.82Win RateUser Study
@@
4.3 Qualitative Evaluation
4.3.1 Analysis of Motion Types
We present qualitative visualizations of our method across different motion types in Fig. 5. Our approach is particularly effective for dance movements, where the inherent compositional structure, characterized by recognizable step combinations, aligns naturally with supermotion segmentation. Similarly, for multi-phase sequential actions such as “walk-kneel-stand”, the supermotion representation intuitively segments actions into meaningful units, preserving structural coherence.
For cyclic locomotion, such as walking and running, our method accurately captures and maintains periodic patterns. The supermotion representation naturally preserves the repetitive nature of these motions while ensuring consistency over extended sequences. Additionally, our approach performs well for large-amplitude motions with stable velocities, such as body twists and bending down. These actions involve smooth velocity transitions, making them particularly suited for Lagrangian motion field modeling. The supermotion representation effectively captures the overall movement trajectory while maintaining natural motion dynamics.
4.3.2 Visualization of Text-Supermotion Alignment
To better understand how our supermotion representation captures semantic meaning from text prompts, we visualize the attention weights between text tokens and motion segments. As shown in Fig. 6, for a sequence involving multiple motion types, “A person waves his arm first, walks forward in a straight line, then turns left,” the attention heatmap reveals distinct peaks corresponding to each action: high attention to “waves” in the initial supermotions, “walks” in the middle segment, and “turns” in the final supermotions. Similarly, for the sequence “A person jumps up, and walks forward,” we observe strong attention weights between the verb “jumps” and the initial motion segments (supermotion 0 to 5), followed by high attention to “walks” in the later segments (supermotion 8 to 14). These visualizations demonstrate that our model effectively aligns specific motion segments with their corresponding action words in the text description, highlighting its ability to capture semantic relationships between language and motion.
4.3.3 Qualitative Comparison
Qualitative results for the text-to-motion and music-to-dance tasks are presented in Fig. 7 and Fig. 8, respectively. Additional examples are provided in the video demonstrations included in the supplemental materials. Compared to baseline methods, our approach generates more diverse and coherent motions over extended durations, demonstrating its effectiveness in long-term motion synthesis.
4.4 User Study